# Will AI Take My Job? Evolving Perceptions of Automation and Labor Risk in Latin America

### v1

@@ Introduction
This paper addresses this gap by analyzing responses from over 48,000 individuals across 16 Latin American countries, drawn from the 2017, 2018, 2020, and 2023 waves of the Latinobarómetro survey (Latinobarómetro Corporation 2017–2023). Using a combination of statistical modeling and latent class analysis, we examine how fear of AI-induced job loss varies across countries, time periods, and social groups. Our analysis highlights education level and political ideology as the most consistent predictors of concern, while also uncovering how these attitudes cluster into distinct societal profiles.
We find that concern about AI-driven job loss peaked in 2018, likely due to changes in question framing and heightened media attention, but remains widespread and unequally distributed. Fear is especially pronounced among individuals with lower educational attainment and among those on the political left. The results underscore the importance of structural conditions and ideological identity in shaping public perceptions of automation in Latin America. In doing so, this study contributes a comparative perspective to a literature still heavily focused on the Global North, offering new insight into how technological anxiety is patterned in emerging economies.
@@
Education. Education is a robust predictor of AI attitudes. Higher education is associated with lower anxiety and greater acceptance (Schiavo, Businaro, and Zancanaro 2024). AI literacy, defined as understanding, using, and critically evaluating AI, is both shaped by and mediates the effect of education on AI acceptance (Carolus et al. 2023; Lund, Mannuru, and Agbaji 2024).
@@
Trust in those developing and governing AI is a strong predictor of public acceptance. While trust in academia and healthcare is relatively high, governments and tech companies are viewed more skeptically (Gillespie, Lockey, and Curtis 2021). This trust gap fuels anxiety: respondents want regulation but question government competence (Gerlich 2023; Montag et al. 2023), and anxiety drops when governance is trusted (Schiavo, Businaro, and Zancanaro 2024; Zhan et al. 2024). A lack of trust heightens concerns about autonomy, transparency, and job loss.
@@
The analysis is based on survey data collected by Latinobarómetro, a public opinion study conducted annually across Latin America. We use the 2017, 2018, 2020, and 2023 editions of the survey, which include responses from 16 countries: Bolivia, Brazil, Chile, Colombia, Costa Rica, Dominican Republic, Ecuador, El Salvador, Guatemala, Honduras, Mexico, Panama, Paraguay, Peru, Uruguay, and Venezuela. Each wave captures perceptions of democracy, governance, economic outlook, and, crucially, emerging technologies like AI.
@@ Results
Figure 3 displays the evolution of estimated coefficients for three key predictors, education (top), political ideology (middle), and city size (bottom), across countries and survey years. Each panel shows country-level estimates for each level of the predictor, with the pooled “FULL” model included for comparison.
@@
For political ideology, coefficients are mostly close to zero, though the Centre category often shows slightly more positive values than Left or Right, suggesting that centrist respondents may express more concern about AI-driven job displacement than ideological extremes.
For city size, respondents in large cities and capitals tend to show more positive coefficients, especially in 2023, indicating greater fear of job loss in urban areas. In contrast, those in mid-sized cities and small towns typically show coefficients closer to zero or negative, pointing to a modest but persistent urban–rural divide.
Full-model estimates in the Appendix confirm these patterns, where education and ideology show some of the largest coefficients across years, underscoring their role as structural drivers of concern over AI-induced job loss.
@@
Unlike other predictors, age is treated as a continuous variable in the ordinal regression models. It is included in approximately 32% of fitted models across countries and years, and when selected, the estimated coefficients are predominantly negative—suggesting that younger individuals tend to express greater concern about AI-driven job loss than older individuals. This negative association is most evident in 2017 and 2023, particularly in Bolivia, Costa Rica, El Salvador, Guatemala, and Peru.
@@
To identify distinct groups of respondents based on their demographic characteristics, economic perceptions, political attitudes, and trust in institutions, we perform a Latent Class Analysis (LCA) using all predictors except country and fear_ai. Although the BIC decreases monotonically with additional classes, likely due to the large sample size, models with more than four classes show diminishing separation and interpretability. We therefore retain a four-class solution as a compromise between BIC and entropy, balancing model fit and classification certainty. The corresponding values of both metrics for 1 to 10 classes are reported in the Appendix.
Each class reflects a distinct combination of respondent characteristics, summarized below. Interpretations are based on the estimated item-response probabilities $\mathbb{P}(X_{j}=x_{j}\mid C=k)$, which represent the likelihood that individuals in latent class $k$ choose a particular category of variable $X_{j}$. These probabilities inform the substantive meaning of each profile, with full tables provided in the Appendix.
@@
  • C3: Moderate Progressives. These respondents tend to be younger, with above-average levels of education and moderate trust in institutions. They report moderate satisfaction with democracy and relatively good expectations for their personal economic future. Their ideological self-placement is more often centrist or left-leaning.
@@ Discussion
Political ideology also plays a prominent and growing role in structuring attitudes toward AI. Respondents identifying with the political left exhibit higher levels of concern about AI-driven job loss than those at the center or right, especially in 2018 and 2023. This mirrors the pattern observed in European studies such as Gerlich (2023), where left-leaning individuals are more likely to perceive AI as a threat to social equality and labor rights, while right-leaning respondents express greater optimism about innovation and market-led adoption. In our data, the ideological gradient appears to widen over time, suggesting that debates over automation are filtered through partisan and ideological lenses.

### v2

@@ Introduction
This paper addresses this gap by analyzing responses from over 48,000 individuals across 16 Latin American countries, drawn from the four most recent waves (2017, 2018, 2020, and 2023) of the Latinobarómetro survey (Latinobarómetro Corporation 2017–2023). Using a combination of statistical modeling and latent class analysis, we examine how fear of AI-induced job loss varies across countries, time periods, and social groups. Our analysis highlights education level and political ideology as the most consistent predictors of concern, while also uncovering how these attitudes cluster into distinct societal profiles.
We find that concern about AI-driven job loss peaked in 2018, likely due to changes in question framing and heightened media attention, but remains widespread and unequally distributed. Fear is especially pronounced among individuals with lower educational attainment. Political ideology also plays a role, with left-leaning respondents expressing greater concern in earlier years, though this pattern becomes less consistent over time and, in some cases, reverses. The results underscore the importance of structural conditions and ideological identity in shaping public perceptions of automation in Latin America. In doing so, this study contributes a comparative perspective to a literature still heavily focused on the Global North, offering new insight into how technological anxiety is patterned in emerging economies.
@@
Education. Education is a robust predictor of AI attitudes. Higher education is associated with lower anxiety and greater acceptance (Schiavo, Businaro, and Zancanaro 2024). AI literacy, defined as understanding, using, and critically evaluating AI, is both shaped by and mediates the effect of education on AI acceptance (Lund, Mannuru, and Agbaji 2024).
@@
Trust in those developing and governing AI is a strong predictor of public acceptance. While trust in academia and healthcare is relatively high, governments and tech companies are viewed more skeptically (Gillespie, Lockey, and Curtis 2021). This trust gap fuels anxiety: respondents want regulation but question government competence (Cremaschi, Lee, and Leonelli 2025; Gerlich 2023; Montag et al. 2023), and anxiety drops when governance is trusted (Schiavo, Businaro, and Zancanaro 2024; Zhan et al. 2024). A lack of trust heightens concerns about autonomy, transparency, and job loss.
@@
The analysis is based on survey data collected by Latinobarómetro, a public opinion study conducted across Latin America. We use the four most recent editions of the survey (2017, 2018, 2020, and 2023), which include responses from 16 countries: Bolivia, Brazil, Chile, Colombia, Costa Rica, Dominican Republic, Ecuador, El Salvador, Guatemala, Honduras, Mexico, Panama, Paraguay, Peru, Uruguay, and Venezuela. Each wave captures perceptions of democracy, governance, economic outlook, and, crucially, emerging technologies like AI.
@@ Results
Figure 3 displays the evolution of estimated coefficients for three key predictors, education (top), political ideology (middle), and city size (bottom), across countries and survey years. Each panel shows country-level estimates for each level of the predictor, with the pooled “FULL” model included for comparison. Because of the model structure and the use of sum-to-zero contrasts, negative coefficient estimates indicate greater fear of AI-driven job loss (i.e., a higher probability of agreeing or strongly agreeing that AI will lead to job displacement).
@@
For political ideology, coefficients are mostly close to zero, though the Centre category consistently shows slightly more positive values than Left or Right, suggesting that centrist respondents tend to express less concern about AI-driven job displacement. That said, the direction of coefficients for Left and Right varies by year: while left-leaning individuals show greater concern in earlier waves (e.g., 2017 and 2018), this pattern is less clear or even reversed by 2023.
For city size, respondents in large cities and capitals tend to show more positive coefficients, especially in 2023, indicating less fear of job loss in urban areas. In contrast, those in mid-sized cities and small towns typically show coefficients closer to zero or negative, pointing to a modest but persistent urban–rural divide.
Full-model estimates (available in the extended version) confirm these patterns, with education and ideology showing some of the largest coefficients across years and supporting their role as structural drivers of concern over AI-induced job loss.
@@
Unlike other predictors, age is treated as a continuous variable in the ordinal regression models. It is included in approximately 32% of fitted models across countries and years, and when selected, the estimated coefficients are predominantly negative, suggesting that younger individuals tend to express greater concern about AI-driven job loss than older individuals. This negative association is most evident in 2017 and 2023, particularly in Bolivia, Costa Rica, El Salvador, and Peru (see the extended version).
@@
To identify distinct groups of respondents based on their demographic characteristics, economic perceptions, political attitudes, and trust in institutions, we perform a Latent Class Analysis (LCA) using all predictors except country and fear_ai. Although the BIC decreases monotonically with additional classes, likely due to the large sample size, models with more than four classes show diminishing separation and interpretability. We therefore retain a four-class solution as a compromise between BIC and entropy, balancing model fit and classification certainty (see the extended version for details).
Each class reflects a distinct combination of respondent characteristics, summarized below. Interpretations are based on the estimated item-response probabilities $\mathbb{P}(X_{j}=x_{j}\mid C=k)$, which represent the likelihood that individuals in latent class $k$ choose a particular category of variable $X_{j}$. These probabilities inform the substantive meaning of each profile and are available in the extended version.
@@
  • C3: Moderate Progressives. These respondents tend to be younger, with above-average levels of education and moderate trust in institutions. They report moderate satisfaction with democracy and relatively good expectations for their personal economic future. Their ideological self-placement is more often centrist.
@@ Discussion
Political ideology plays a nuanced role in shaping attitudes toward AI. Across most years, respondents identifying with the political centre tend to express slightly less concern about AI-driven job loss compared to those on the ideological left or right. While earlier waves (e.g., 2017 and 2018) show somewhat greater fear among left-leaning individuals, this pattern becomes less consistent over time, with right-leaning respondents exhibiting greater concern in some countries by 2023. These findings partially mirror results from European studies such as Gerlich (2023), which highlight ideological divides in perceptions of automation, though in our data the gradient is modest and appears to shift over time rather than follow a fixed partisan pattern.
@@ Appendix A Additional Results
Country 2017 2018 2020 2023
Bolivia -0.0072 -0.0089 -0.0140
Brazil -0.0108
Chile -0.0060
Colombia -0.0138
Costa Rica -0.0151 -0.0067
Dominican Republic
Ecuador
El Salvador -0.0116 -0.0103
FULL -0.0031 -0.0023
Guatemala -0.0138 0.0148
Honduras
Mexico 0.0127
Panama -0.0095
Paraguay 0.0126
Peru -0.0121 -0.0103
Uruguay 0.0229
Venezuela -0.0055
Table 6: Estimated coefficients for age in country-specific and pooled (FULL) models across survey years. Blank cells indicate that age was not included in the model for that country-year.